# A Large Deviations Perspective on Policy Gradient Algorithms

### v2

@@ Abstract
$\begin{split}&V^{\theta_{t+1}}_{\tau}(\rho)-V_{\tau}^{\theta^{\star}}(\rho)\\ &\leq(1-\mu\eta_{t}+\mu\eta_{t}^{2}L_{1})(V^{\theta_{t}}_{\tau}(\rho)-V_{\tau}^{\theta^{\star}}(\rho))+\eta_{t}\langle g(\theta_{t}),Z_{t}\rangle+{\eta_{t}^{2}L_{1}}\|Z_{t}\|_{2}^{2}.\end{split}$ (3.1)
Also, observe that since [^2] To show that $\mu\leq L_{1}$, one exploits that $L_{1}$-smoothness implies Lipschitz continuity of the gradient and recalls that the PL condition implies a quadratic growth condition [KNS16, App. A]. $\eta_{t}\leq 1/L_{1}$, $\mu\leq L_{1}$ and $\eta_{t},L_{1}>0$, we have $\mu\eta_{t}(\eta_{t}L_{1}-1)\in(-1,0]$ and thus $(1-\mu\eta_{t}+\mu\eta_{t}^{2}L_{1})\in(0,1]$ for $t=1,\dots,T$. Next, by substituting the points $\theta^{\prime}=\theta-1/L_{1}\cdot g(\theta)$ and $\theta$ into (2.5), we obtain $V^{\theta-1/L_{1}\cdot g(\theta)}_{\tau}(\rho)\leq V^{\theta}_{\tau}(\rho)-\tfrac{1}{2}L_{1}^{-1}\|g(\theta)\|_{2}^{2}$. Using the optimality of $\theta^{\star}$, we may then conclude that $V_{\tau}^{\theta^{\star}}(\rho)\leq V^{\theta-1/L_{1}\cdot g(\theta)}_{\tau}(\rho)\leq V^{\theta}_{\tau}(\rho)-\tfrac{1}{2}L_{1}^{-1}\|g(\theta)\|_{2}^{2}$ which implies that
$\|g(\theta)\|_{2}^{2}$ $\leq 2L_{1}(V^{\theta}_{\tau}(\rho)-V^{\theta^{\star}}_{\tau}(\rho))\quad\forall\theta\in\mathbb{R}^{d}.$ (3.2)
@@
Suppose Assumptions 2.12.2, and 2.3 hold, and choose $T>1$, $\epsilon\in(0,1)$. Then, there is a universal constant $C>0$ such that the following hold with probability at least $61-\epsilon/$ 6. Let $C_{\mathsf{M}}=(\sigma\sqrt{|\mathcal{S}||\mathcal{A}|}C)^{2}$ where $\sigma$ is as in (2.4). and set $\eta>0$ such that $(\mu\eta-1)>\sigma^{2}/C_{\mathsf{M}}$. Let $\eta_{t}=\eta/(t+t_{0}+1)$ for all $t=1,\ldots,T$ and choose $t_{0}$ and $K$ such that
$t_{0}\geq$ $\max\left\{\frac{\eta^{2}L_{1}}{(\mu\eta-1)-B_{0}C_{0}\eta^{2}}-1,L_{1}\eta-2,\sqrt{\frac{3\sigma^{2}}{2\epsilon\Delta}}\right\},$
$K\geq$ $\max_{t=1,\dots,T}\left\{B_{0}^{-1},(t_{0}+1)\big{(}V^{\theta_{1}}_{\tau}(\rho)-V_{\tau}^{\theta^{\star}}(\rho)\big{)},\frac{2c_{t}C_{\mathsf{M}}}{1-(a_{t}+B_{0}C_{0}b_{t}^{2})}\right\},$ (3.3)
@@
[…]
Then, for any $\delta\geq 0$ and $t=1,\dots,T$ we have that
$\mathbb{P}\left(\big{(}V^{\theta_{t+1}}_{\tau}(\rho)-V_{\tau}^{\theta^{\star}}(\rho)\big{)}\geq\delta\right)\leq e^{1-(t+t_{0}+1)\delta/K}.$
The proof is inspired by [BJK23], with the difference being that they work with strongly convex objective functions. Unfortunately, $V^{\theta}_{\tau}(\rho)$ fails to be convex in $\theta.$
Proof.
Define an auxiliary process $Y_{t}=(t+t_{0})(V^{\theta_{t}}_{\tau}(\rho)-V_{\tau}^{\theta^{\star}}(\rho)).$ Then, the recursion (3.1) implies that $Y_{t+1}\leq a_{t}Y_{t}+b_{t}\sqrt{t+t_{0}}\langle g(\theta_{t}),Z_{t}\rangle+c_{t}\|Z_{t}\|^{2},$ where $a_{t}$, $b_{t}$ and $c_{t}$ are as in (3.4). Defining the MGFs associated with $Y_{t}$ as $\Phi_{t}(\nu)=\mathbb{E}\left[\exp\left(\nu Y_{t}\right)\right]$ and $\Phi_{t+1|t}\left(\nu;\theta_{t}\right)=\mathbb{E}\left[\exp\left(\nu Y_{t+1}\right)\mid\theta_{t}\right]$ for $\nu\in\mathbb{R}$, the recursion (3.1) further implies that
$\Phi_{t+1|t}\left(\nu;\theta_{t}\right)$
$\!\leq\exp\left(a_{t}\nu Y_{t}\right)\mathbb{E}\left[\exp\left(2b_{t}\nu\sqrt{t+t_{0}}\langle g(\theta_{t}),Z_{t}\rangle\right)\mid\theta_{t}\right]^{1/2}\mathbb{E}\!\left[\exp\big{(}2c_{t}\nu\left\|Z_{t}\right\|_{2}^{2}\big{)}\mid\theta_{t}\right]^{1/2}$
@@
where the first inequality follows from Hölder’s inequality, the second inequality follows from (2.4) and the last inequality follows from (3.2). Recall that $d=\mathrm{dim}(Z_{t})$ and define $B_{0}=1/(2\eta^{2}L_{1}(\sigma\sqrt{d}C)^{2})$. Then, there is a constant $C>0$ such that for all $\nu\in[0,B_{0}]$, we have by monotonicity of $c_{t}$, Hölder’s inequality for $p=1/(\nu 2\eta^{2}L_{1}(\sigma\sqrt{d}C)^{2})\geq 1$ and [Jin+19, Lem. 2] that $M(2c_{t}\nu;\theta)\leq\mathrm{exp}(2c_{t}\nu(\sigma\sqrt{d}C)^{2})$. Recall that $(\sigma\sqrt{d}C)^{2}=C_{\mathsf{M}}$ because $d=|\mathcal{S}||\mathcal{A}|$. Moreover, as $\mathrm{exp}(x)\geq\mathrm{exp}(x)^{1/2}$ for $x\geq 0$, we can simplify the above inequality to
$\Phi_{t+1|t}\left(\nu;\theta_{t}\right)$ $\leq\exp\left(a_{t}\nu Y_{t}\right)\exp\left(2L_{1}\sigma^{2}b_{t}^{2}\nu^{2}Y_{t}\right)\mathrm{exp}(2c_{t}\nu C_{\mathsf{M}})$ (3.5)
@@
for $C_{0}=2L_{1}\sigma^{2}$. Taking expectations on both sides of (3.5) yields
$\Phi_{t+1}(\nu)\leq\Phi_{t}\left(\nu(a_{t}+B_{0}C_{0}b_{t}^{2})\right)\mathrm{exp}(2c_{t}\nu C_{\mathsf{M}})\quad\forall\nu\leq B_{0}.$ (3.6)
@@
$a_{t}=1-\frac{\mu\eta-1}{t+t_{0}}\left(1-\frac{\mu\eta^{2}L_{1}}{(\mu\eta-1)(t+t_{0}+1)}\right),$ (3.7)
where we exploit the identity $1+1/s=(s+1)/s$ for $s\neq 0$. We have that $(\mu\eta-1)>0$. Now since $t_{0}\geq L_{1}\eta-2$ by assumption, we have $\eta_{t}\leq 1/L_{1}$. Recall as well the fact that $\mu\leq L_{1}$, which leads to $\eta_{t}\leq 1/\mu$ and thus $\mu\eta_{t}\leq 1$, that is, $\mu\eta/(t+t_{0}+1)\leq 1$. Exactly this implies that $(\mu\eta-1)/(t+t_{0})\in(0,1]$. This covers the first non-trivial fraction of (3.7). Now for the second part, we have the implication that
@@
from where we can conclude that $a_{t}\in[0,1)$ for any finite $t\geq 1$. However, we will need $(a_{t}+B_{0}C_{0}b_{t}^{2})<1$. As $b_{t}$ decays with $t_{0}$, this can be achieved by selecting a sufficiently large $t_{0}$. Explicitly, let $\mu\eta=1+\varepsilon$ for some $\varepsilon>0$. By the definition of $b_{t}$ and $a_{t}$ as in (3.7) we then obtain
$a_{t}+B_{0}C_{0}b_{t}^{2}=1-\frac{\varepsilon}{t+t_{0}}\left(1-\frac{B_{0}C_{0}\eta^{2}}{\varepsilon}-\frac{\mu\eta^{2}L_{1}}{\varepsilon(t+t_{0}+1)}\right).$ (3.8)
By assumption, we have $\mu\eta-1>\sigma^{2}/C_{\mathsf{M}}.$ This implies that $\varepsilon>B_{0}C_{0}\eta^{2}.$ It then follows that $a_{t}+B_{0}C_{0}b_{t}^{2}\in[0,1)$ for any finite $t\geq 1$ and $t_{0}+1\geq\mu\eta^{2}L_{1}/(\varepsilon-B_{0}C_{0}\eta^{2}).$ Recall also that $t_{0}\geq L_{1}\eta-2$ by assumption. Thus, we have $\eta_{t}\leq 1/L_{1}$. Assumption 2.1 allows then for the following inductive procedure. Let $\nu\in[0,1/K]$ for $K$ as in (3.3) then, for $t=1$, we have $\Phi_{1}(\nu)\leq\mathrm{exp}(\nu K)$ since $K\geq t_{0}\left(V^{\theta_{1}}_{\tau}(\rho)-V_{\tau}^{\theta^{\star}}(\rho)\right)$. Now suppose $\Phi_{t}(\nu)\leq\mathrm{exp}(\nu K)$ for some arbitrary $t\geq 1$, then, we have by the inductive assumption that $\Phi_{t+1}\leq\mathrm{exp}(2c_{t}C_{\mathsf{M}}\nu)\mathrm{exp}(\nu(a_{t}+B_{0}C_{0}b_{t}^{2})K).$ However, since for $t=1,\dots,T$ we also have that $K\geq 2c_{t}C_{\mathsf{M}}/(1-(a_{t}+B_{0}C_{0}b_{t}^{2})),$ which is well-defined since $(a_{t}+B_{0}C_{0}b_{t}^{2})<1$, by our selection of $t_{0}$ and $\eta$, it follows that $\Phi_{t+1}(\nu)\leq\mathrm{exp}(\nu K)$ for $t=1,\dots,T$. By [Har+19, Claim A.7], if a random variable $X\in\mathbb{R}$ satisfies $\mathbb{E}[\mathrm{exp}(\lambda X)]\leq C_{1}\mathrm{exp}(\lambda C_{2})$ for all $\lambda\leq 1/C_{2}$ and some universal constants $C_{1}$ and $C_{2}$, then $\mathbb{P}(X\geq C_{2}\log(1/\delta))\leq C_{1}e\delta$. Hence, after applying [Har+19, Claim A.7] one finds that $\mathbb{P}(Y_{t}\geq C_{3})\leq\mathrm{exp}(1-C_{3}/K),$ for any choice of $C_{3}\geq 0$ and $t=1,\dots,T$. The proof is concluded by substituting the expression for $Y_{t}$ into the previous inequality and a union bound. ∎
We note that by the expressions for $(a_{t},b_{t},c_{t})$ and (3.8) it follows that $\lim_{t\to+\infty}2c_{t}C_{\mathsf{M}}/$ $(1-(a_{t}+B_{0}C_{0}b_{t}^{2}))$ is bounded, that is, one can freely select $T$, which we will exploit later in Theorem 4.2.
@@
Assumption 4.1 (Conditional LMGF regularity).
@@
We refer to [BJK23, p. 5] for a discussion of this assumption. For instance, when the conditional distribution of $Z_{t}$ given $\theta_{t}=\theta$ is Gaussian, and its covariance matrix is Lipschitz continuous in $\theta$, then Assumption 4.1 holds. In particular, if the covariance is independent of $\theta$, this trivially holds true. Next, we define the MGF of $\theta_{t}$ as $\mathbb{R}^{d}\ni\lambda\mapsto\Gamma_{t}(\lambda):=\mathbb{E}\left[\mathrm{exp}(\langle\lambda,\theta_{t}-\theta^{\star}\rangle)\right]$ and similarly the LMGF as $\mathbb{R}^{d}\ni\lambda\mapsto\log\Gamma_{t}(\lambda)$. Now we are equipped to provide our main technical result, leading directly to a large deviation principle (LDP) upper bound. Note that $\Gamma_{t}(\lambda)$ depends on the optimal solution $\theta^{\star}$ which is fixed and unknown.
@@
Suppose that Assumptions 2.12.2,  2.3 and 4.1 hold and set $\eta_{t}=\eta/(t+t_{0}+1)$ as in Lemma 3.2. Then, with probability at least $61-\epsilon/$ 6 we have that
$\limsup_{t\to+\infty}\frac{1}{t}\log\Gamma_{t}(t\lambda)\leq r(\lambda)+\int_{0}^{1}\Lambda\left(\eta QD(x)Q^{\top}\lambda;\theta^{\star}\right)\mathrm{d}x=:\Psi(\lambda),$ (4.1)
@@
To prove Theorem 4.2, we will largely follow the proof strategy of [BJK23, Lem. 6], yet, with technical deviations regarding indexing of the sequences. In addition, we cannot appeal to the strong convexity exploited in [BJK23].
Proof.
We provide a brief overview of the proof. Step 0 constructs a deterministic sequence as a basis for deriving recursions of MGF. Step 1 defines a ball around the optimal parameter and split the cases depending on whether $\theta_{\ell}$ resides within the ball. Step 2-4 study individual cases leading to their respective recursive relation. Finally, Step 5 summarizes the previously derived recursions, takes the limit and concludes. Step 0, sequence. Fix some vector $\lambda\in\mathbb{R}^{d}$, integer $t\geq 1$ and define the sequence of vectors $\zeta_{\ell}=B_{t,\ell}\zeta_{t}$ via $B_{t,\ell}=(I-\eta_{\ell}H(\theta^{\star}))\cdot...\cdot(I-\eta_{t-1}H(\theta^{\star}))$ and $\zeta_{t}=t\lambda$, with $1\leq\ell<t$. As $\eta_{t}\leq 1/L_{1}$ we have that $\|I-\eta_{t}H(\theta^{\star})\|_{2}\leq 1-\eta_{t}\lambda_{\mathrm{min}}(H(\theta^{\star}))$. Hence, using [BJK23, Lem. 2] and exploiting ${\eta\lambda_{\mathrm{min}}(H(\theta^{\star}))}>1$i.e., recall $\eta\mu>1$ from Lemma 3.2, one can show that
@@
Step 1, conditional MGF. Let $\mu_{\ell}$ and $\nu_{\ell}$ be Borel measures induced by $\theta_{\ell}$ and $\|\theta_{\ell}-\theta^{\star}\|_{2}$, respectively. Then, one readily shows that $\Gamma_{\ell+1}(\zeta_{\ell})=\textstyle\int_{\mathbb{R}^{d}}\Gamma_{\ell+1|\ell}(\zeta_{\ell};\theta)\mu_{\ell}(\mathrm{d}\theta).$ Next, define the sets $B_{\ell}(\theta^{\star},\delta)=\{\theta_{\ell}:\|\theta_{\ell}-\theta^{\star}\|_{2}\leq\delta\}$. Now we can construct the decomposition $\Gamma_{\ell+1}(\zeta_{\ell})=\Gamma_{\ell+1|B_{\ell}(\theta^{\star},\delta)}(\zeta_{\ell})+\Gamma_{\ell+1|B^{c}_{\ell}(\theta^{\star},\delta)}(\zeta_{\ell})$, which we study separately. Step 2, $\theta\in B_{\ell}(\theta^{\star},\delta)$. We have that
$\Gamma_{\ell+1|\ell}(\zeta_{\ell};\theta)=\mathrm{exp}\left(\Lambda(\eta_{\ell}\zeta_{\ell};\theta)+\langle\zeta_{\ell},\theta-\eta_{\ell}g(\theta)-\theta^{\star}\rangle\right).$
Let $H(\theta)$ denote the Hessian of $V^{\pi_{\theta}}_{\tau}(\rho)$ at $\theta$ and define the residual term $h(\theta)=g(\theta)-H(\theta^{\star})(\theta-\theta^{\star})$. Recall Step 0, Assumption 4.1 and define the largest residual term $\bar{h}(\delta)=\sup_{\theta\in B_{\ell}(\theta^{\star},\delta)}\|h(\theta)\|_{2}$, it follows that
@@
$\Gamma_{\ell+1|B_{\ell}(\theta^{\star},\delta)}(\zeta_{\ell})\leq\mathrm{exp}\left(\Lambda(\eta_{\ell}\zeta_{\ell};\theta^{\star})+r_{0}(\lambda,\delta)\right)\Gamma_{\ell}(\zeta_{\ell-1}).$ (4.2)

### v3

@@ Abstract
$V^{\theta_{t+1}}_{\tau}(\rho)-V_{\tau}^{\theta^{\star}}(\rho)\leq(1-\mu\eta_{t}+\mu\eta_{t}^{2}L_{1})(V^{\theta_{t}}_{\tau}(\rho)-V_{\tau}^{\theta^{\star}}(\rho))+\eta_{t}\langle g(\theta_{t}),Z_{t}\rangle+{\eta_{t}^{2}L_{1}}\|Z_{t}\|_{2}^{2}.$ (7)
Also, observe that since [^2] To show that $\mu\leq L_{1}$, one exploits that $L_{1}$-smoothness implies Lipschitz continuity of the gradient and recalls that the PL condition implies a quadratic growth condition (Karimi et al., 2016, App. A). $\eta_{t}\leq 1/L_{1}$, $\mu\leq L_{1}$ and $\eta_{t},L_{1}>0$, we have $\mu\eta_{t}(\eta_{t}L_{1}-1)\in(-1,0]$ and thus $(1-\mu\eta_{t}+\mu\eta_{t}^{2}L_{1})\in(0,1]$ for $t=1,\dots,T$. Next, by substituting the points $\theta^{\prime}=\theta-1/L_{1}\cdot g(\theta)$ and $\theta$ into (6), we obtain $V^{\theta-1/L_{1}\cdot g(\theta)}_{\tau}(\rho)\leq V^{\theta}_{\tau}(\rho)-\tfrac{1}{2}L_{1}^{-1}\|g(\theta)\|_{2}^{2}$. Using the optimality of $\theta^{\star}$, we may then conclude that $V_{\tau}^{\theta^{\star}}(\rho)\leq V^{\theta-1/L_{1}\cdot g(\theta)}_{\tau}(\rho)\leq V^{\theta}_{\tau}(\rho)-\tfrac{1}{2}L_{1}^{-1}\|g(\theta)\|_{2}^{2}$ which implies that
$\|g(\theta)\|_{2}^{2}$ $\leq 2L_{1}(V^{\theta}_{\tau}(\rho)-V^{\theta^{\star}}_{\tau}(\rho))\quad\forall\theta\in\mathbb{R}^{d}.$ (8)
@@
Suppose Assumptions 12, and 3 hold, and choose $T>1$, $\epsilon\in(0,1)$. Then, there is a universal constant $C>0$ such that the following event holds with probability at least $61-\epsilon/$ 6. Let $C_{\mathsf{M}}=(\sigma\sqrt{|\mathcal{S}||\mathcal{A}|}C)^{2}$ where $\sigma$ is as in (5), and set $\eta>0$ such that $(\mu\eta-1)>\sigma^{2}/C_{\mathsf{M}}$. Let $\eta_{t}=\eta/(t+t_{0}+1)$ for all $t=1,\ldots,T$ and choose $t_{0}$ and $K$ such that
$t_{0}\geq$ $\max\left\{\frac{\eta^{2}L_{1}}{(\mu\eta-1)-B_{0}C_{0}\eta^{2}}-1,L_{1}\eta-2,\sqrt{\frac{3\sigma^{2}}{2\epsilon\Delta}}-1\right\},$
$K\geq$ $\max_{t=1,\dots,T}\left\{B_{0}^{-1},(t_{0}+1)\big{(}V^{\theta_{1}}_{\tau}(\rho)-V_{\tau}^{\theta^{\star}}(\rho)\big{)},\frac{2c_{t}C_{\mathsf{M}}}{1-(a_{t}+B_{0}C_{0}b_{t}^{2})}\right\},$ (9)
@@
$a_{t}=\frac{t+t_{0}+1}{t+t_{0}}\left(1-\mu\eta_{t}+\mu\eta_{t}^{2}L_{1}\right),\ b_{t}=\frac{\eta}{\sqrt{t+t_{0}}},\ \text{and}\ c_{t}=\frac{\eta^{2}L_{1}}{t+t_{0}+1}.$ (10)
Then, for any $\delta\geq 0$ and $t=1,\dots,T$ we have that $\mathbb{P}\left(\big{(}V^{\theta_{t+1}}_{\tau}(\rho)-V_{\tau}^{\theta^{\star}}(\rho)\big{)}\geq\delta\right)\leq e^{1-(t+t_{0}+1)\delta/K}.$
The proof is inspired by (Bajovic et al., 2023), with the difference being that they work with strongly convex objective functions. Unfortunately, $V^{\theta}_{\tau}(\rho)$ fails to be convex in $\theta.$
Proof 3.4.
Define an auxiliary process $Y_{t}=(t+t_{0})(V^{\theta_{t}}_{\tau}(\rho)-V_{\tau}^{\theta^{\star}}(\rho)).$ Then, the recursion (7) implies that $Y_{t+1}\leq a_{t}Y_{t}+b_{t}\sqrt{t+t_{0}}\langle g(\theta_{t}),Z_{t}\rangle+c_{t}\|Z_{t}\|^{2},$ where $a_{t}$, $b_{t}$ and $c_{t}$ are as in (10). Defining the MGFs associated with $Y_{t}$ as $\Phi_{t}(\nu)=\mathbb{E}\left[\exp\left(\nu Y_{t}\right)\right]$ and $\Phi_{t+1|t}\left(\nu;\theta_{t}\right)=\mathbb{E}\left[\exp\left(\nu Y_{t+1}\right)\mid\theta_{t}\right]$ for $\nu\in\mathbb{R}$, the recursion (7) further implies that
$\Phi_{t+1|t}\left(\nu;\theta_{t}\right)$ $\!\leq\exp\left(a_{t}\nu Y_{t}\right)\mathbb{E}\left[\exp\left(2b_{t}\nu\sqrt{t+t_{0}}\langle g(\theta_{t}),Z_{t}\rangle\right)\mid\theta_{t}\right]^{1/2}\mathbb{E}\!\left[\exp\big{(}2c_{t}\nu\left\|Z_{t}\right\|_{2}^{2}\big{)}\mid\theta_{t}\right]^{1/2}$
@@
where the first inequality follows from Hölder’s inequality, the second inequality follows from (5) and the last inequality follows from (8). Recall that $d=\mathrm{dim}(Z_{t})$ and define $B_{0}=1/(2\eta^{2}L_{1}(\sigma\sqrt{d}C)^{2})$. Then, there is a constant $C>0$ such that for all $\nu\in[0,B_{0}]$, we have by monotonicity of $c_{t}$, Hölder’s inequality for $p=1/(\nu 2\eta^{2}L_{1}(\sigma\sqrt{d}C)^{2})\geq 1$ and (Jin et al., 2019, Lem. 2) that $M(2c_{t}\nu;\theta)\leq\mathrm{exp}(2c_{t}\nu(\sigma\sqrt{d}C)^{2})$. Recall that $(\sigma\sqrt{d}C)^{2}=C_{\mathsf{M}}$ because $d=|\mathcal{S}||\mathcal{A}|$. Moreover, as $\mathrm{exp}(x)\geq\mathrm{exp}(x)^{1/2}$ for $x\geq 0$, we can simplify the above inequality to
$\Phi_{t+1|t}\left(\nu;\theta_{t}\right)$ $\leq\exp\left(a_{t}\nu Y_{t}\right)\exp\left(2L_{1}\sigma^{2}b_{t}^{2}\nu^{2}Y_{t}\right)\mathrm{exp}(2c_{t}\nu C_{\mathsf{M}})$ (11)
@@
for $C_{0}=2L_{1}\sigma^{2}$. Taking expectations on both sides of (11) yields
$\Phi_{t+1}(\nu)\leq\Phi_{t}\left(\nu(a_{t}+B_{0}C_{0}b_{t}^{2})\right)\mathrm{exp}(2c_{t}\nu C_{\mathsf{M}})\quad\forall\nu\leq B_{0}.$ (12)
@@
$a_{t}=1-\frac{\mu\eta-1}{t+t_{0}}\left(1-\frac{\mu\eta^{2}L_{1}}{(\mu\eta-1)(t+t_{0}+1)}\right),$ (13)
where we exploit the identity $1+1/s=(s+1)/s$ for $s\neq 0$. We have that $(\mu\eta-1)>0$. Now since $t_{0}\geq L_{1}\eta-2$ by assumption, we have $\eta_{t}\leq 1/L_{1}$. Recall as well the fact that $\mu\leq L_{1}$, which leads to $\eta_{t}\leq 1/\mu$ and thus $\mu\eta_{t}\leq 1$, that is, $\mu\eta/(t+t_{0}+1)\leq 1$. Exactly this implies that $(\mu\eta-1)/(t+t_{0})\in(0,1]$. This covers the first non-trivial fraction of (13). Now for the second part, we have the implication that
@@
from where we can conclude that $a_{t}\in[0,1)$ for any finite $t\geq 1$. However, we will need $(a_{t}+B_{0}C_{0}b_{t}^{2})<1$. As $b_{t}$ decays with $t_{0}$, this can be achieved by selecting a sufficiently large $t_{0}$. Explicitly, let $\mu\eta=1+\varepsilon$ for some $\varepsilon>0$. By the definition of $b_{t}$ and $a_{t}$ as in (13) we then obtain
$a_{t}+B_{0}C_{0}b_{t}^{2}=1-\frac{\varepsilon}{t+t_{0}}\left(1-\frac{B_{0}C_{0}\eta^{2}}{\varepsilon}-\frac{\mu\eta^{2}L_{1}}{\varepsilon(t+t_{0}+1)}\right).$ (14)
By assumption, we have $\mu\eta-1>\sigma^{2}/C_{\mathsf{M}}.$ This implies that $\varepsilon>B_{0}C_{0}\eta^{2}.$ It then follows that $a_{t}+B_{0}C_{0}b_{t}^{2}\in[0,1)$ for any finite $t\geq 1$ and $t_{0}+1\geq\mu\eta^{2}L_{1}/(\varepsilon-B_{0}C_{0}\eta^{2}).$ Recall also that $t_{0}\geq L_{1}\eta-2$ by assumption. Thus, we have $\eta_{t}\leq 1/L_{1}$. Assumption 1 allows then for the following inductive procedure. Let $\nu\in[0,1/K]$ for $K$ as in (9) then, for $t=1$, we have $\Phi_{1}(\nu)\leq\mathrm{exp}(\nu K)$ since $K\geq t_{0}\left(V^{\theta_{1}}_{\tau}(\rho)-V_{\tau}^{\theta^{\star}}(\rho)\right)$. Now suppose $\Phi_{t}(\nu)\leq\mathrm{exp}(\nu K)$ for some arbitrary $t\geq 1$, then, we have by the inductive assumption that $\Phi_{t+1}\leq\mathrm{exp}(2c_{t}C_{\mathsf{M}}\nu)\mathrm{exp}(\nu(a_{t}+B_{0}C_{0}b_{t}^{2})K).$ However, since for $t=1,\dots,T$ we also have that $K\geq 2c_{t}C_{\mathsf{M}}/(1-(a_{t}+B_{0}C_{0}b_{t}^{2})),$ which is well-defined since $(a_{t}+B_{0}C_{0}b_{t}^{2})<1$, by our selection of $t_{0}$ and $\eta$, it follows that $\Phi_{t+1}(\nu)\leq\mathrm{exp}(\nu K)$ for $t=1,\dots,T$. By (Harvey et al., 2019, Claim A.7), if a random variable $X\in\mathbb{R}$ satisfies $\mathbb{E}[\mathrm{exp}(\lambda X)]\leq C_{1}\mathrm{exp}(\lambda C_{2})$ for all $\lambda\leq 1/C_{2}$ and some universal constants $C_{1}$ and $C_{2}$, then $\mathbb{P}(X\geq C_{2}\log(1/\delta))\leq C_{1}e\delta$. Hence, after applying (Harvey et al., 2019, Claim A.7) one finds that $\mathbb{P}(Y_{t}\geq C_{3})\leq\mathrm{exp}(1-C_{3}/K),$ for any choice of $C_{3}\geq 0$ and $t=1,\dots,T$. The proof is concluded by substituting the expression for $Y_{t}$ into the previous inequality and a union bound.
We note that by the expressions for $(a_{t},b_{t},c_{t})$ and (14) it follows that $\lim_{t\to+\infty}2c_{t}C_{\mathsf{M}}/(1-(a_{t}+B_{0}C_{0}b_{t}^{2}))$ is bounded, that is, one can freely select $T$, which we will exploit later in Theorem 4.1.
@@
Assumption 4 (Conditional LMGF regularity)
@@
We refer to (Bajovic et al., 2023, p. 5) for a discussion of this assumption. For instance, when the conditional distribution of $Z_{t}$ given $\theta_{t}=\theta$ is Gaussian, and its covariance matrix is Lipschitz continuous in $\theta$, then Assumption 4 holds. In particular, if the covariance is independent of $\theta$, this trivially holds true. Next, we define the MGF of $\theta_{t}$ as $\mathbb{R}^{d}\ni\lambda\mapsto\Gamma_{t}(\lambda):=\mathbb{E}\left[\mathrm{exp}(\langle\lambda,\theta_{t}-\theta^{\star}\rangle)\right]$ and similarly the LMGF as $\mathbb{R}^{d}\ni\lambda\mapsto\log\Gamma_{t}(\lambda)$. Now we are equipped to provide our main technical result, leading directly to a large deviation principle (LDP) upper bound. Note that $\Gamma_{t}(\lambda)$ depends on the optimal solution $\theta^{\star}$ which is fixed and unknown.
@@
Suppose that Assumptions 12,  3 and 4 hold and set $\eta_{t}=\eta/(t+t_{0}+1)$ as in Lemma 3.3. Then, conditioning on the event that $\mu\!=\inf_{t\geq 1}\mu(\theta_{t})\!>\!0$, we have that
$\limsup_{t\to+\infty}\frac{1}{t}\log\Gamma_{t}(t\lambda)\leq r(\lambda)+\int_{0}^{1}\Lambda\left(\eta QD(x)Q^{\top}\lambda;\theta^{\star}\right)\mathrm{d}x=:\Psi(\lambda),$ (15)
@@
To prove Theorem 4.1, we will largely follow the proof strategy of (Bajovic et al., 2023, Lem. 6), yet, with technical deviations regarding indexing of the sequences. In addition, we cannot appeal to the strong convexity exploited in (Bajovic et al., 2023).
Proof 4.2.
We provide a brief overview of the proof. Step 0 constructs a deterministic sequence as a basis for deriving recursions of MGF. Step 1 defines a ball around the optimal parameter and split the cases depending on whether $\theta_{\ell}$ resides within the ball. Step 2-4 study individual cases leading to their respective recursive relation. Finally, Step 5 summarizes the previously derived recursions, takes the limit and concludes. Step 0, sequence. Fix some vector $\lambda\in\mathbb{R}^{d}$, integer $t\geq 1$ and define the sequence of vectors $\zeta_{\ell}=B_{t,\ell}\zeta_{t}$ via $B_{t,\ell}=(I-\eta_{\ell}H(\theta^{\star}))\cdot...\cdot(I-\eta_{t-1}H(\theta^{\star}))$ and $\zeta_{t}=t\lambda$, with $1\leq\ell<t$. As $\eta_{t}\leq 1/L_{1}$ we have that $\|I-\eta_{t}H(\theta^{\star})\|_{2}\leq 1-\eta_{t}\lambda_{\mathrm{min}}(H(\theta^{\star}))$. Hence, using (Bajovic et al., 2023, Lem. 2) and exploiting ${\eta\lambda_{\mathrm{min}}(H(\theta^{\star}))}>1$, i.e., recall $\eta\mu>1$ from Lemma 3.3, one can show that
@@
Step 1, conditional MGF. Let $\mu_{\ell}$ and $\nu_{\ell}$ be Borel measures induced by $\theta_{\ell}$ and $\|\theta_{\ell}-\theta^{\star}\|_{2}$, respectively. Then, one readily shows that $\Gamma_{\ell+1}(\zeta_{\ell})=\textstyle\int_{\mathbb{R}^{d}}\Gamma_{\ell+1|\ell}(\zeta_{\ell};\theta)\mu_{\ell}(\mathrm{d}\theta).$ Next, define the sets $B_{\ell}(\theta^{\star},\delta)=\{\theta_{\ell}:\|\theta_{\ell}-\theta^{\star}\|_{2}\leq\delta\}$. Now we can construct the decomposition $\Gamma_{\ell+1}(\zeta_{\ell})=\Gamma_{\ell+1|B_{\ell}(\theta^{\star},\delta)}(\zeta_{\ell})+\Gamma_{\ell+1|B^{c}_{\ell}(\theta^{\star},\delta)}(\zeta_{\ell})$, which we study separately. Step 2, $\theta\in B_{\ell}(\theta^{\star},\delta)$. We have that $\Gamma_{\ell+1|\ell}(\zeta_{\ell};\theta)=\mathrm{exp}\left(\Lambda(\eta_{\ell}\zeta_{\ell};\theta)+\langle\zeta_{\ell},\theta-\eta_{\ell}g(\theta)-\theta^{\star}\rangle\right).$ Let $H(\theta)$ denote the Hessian of $V^{\pi_{\theta}}_{\tau}(\rho)$ at $\theta$ and define the residual term $h(\theta)=g(\theta)-H(\theta^{\star})(\theta-\theta^{\star})$. Recall Step 0, Assumption 4 and define the largest residual term $\bar{h}(\delta)=\sup_{\theta\in B_{\ell}(\theta^{\star},\delta)}\|h(\theta)\|_{2}$, it follows that
@@
$\Gamma_{\ell+1|B_{\ell}(\theta^{\star},\delta)}(\zeta_{\ell})\leq\mathrm{exp}\left(\Lambda(\eta_{\ell}\zeta_{\ell};\theta^{\star})+r_{0}(\lambda,\delta)\right)\Gamma_{\ell}(\zeta_{\ell-1}).$ (16)